A solution is made by dissolving 1.25 g of C2H5OH in 11.6 g of H2O. Calculate the mass percent of C2H5OH, the mass percent of H2O, molality of C2H5OH.
1
Expert's answer
2021-02-11T04:47:18-0500

Mass of C2H5OH = 1.25 g

Mass of H2O = 11.6 g

Mass percent of C2H5 OH

= 1.25 × 100 / 1.25 + 11.6

= 125 / 12.86

= 9.72 .


Mass percent of H2O = 100 - 9.72 = 90.28 %.


Molality

= moles of ethyl Alcohol/mass of solvent in kg


= 1.25 / 46 / 11.6 / 1000

= 1.25 × 1000 / 11.6 × 46

= 2.342578 mol / kg .